# LLM & HPC:Benchmarking DeepSeek’s Performance in High-Performance Computing Tasks

### v1

@@ 1 Introduction
Large Language Models (LLMs) have rapidly advanced, potentially approaching Artificial General Intelligence (AGI). Trained on vast textual data, LLMs have shown remarkable capabilities in natural language processing and visualization tasks [1, 12, 22]. This trend has accelerated with the release of the DeepSeek models by the Chinese company DeepSeek [5]. Their DeepSeek-V3 [17, 14] was trained in two months at a cost of \$5.6 million, significantly cheaper than similar models. On January 20th, they launched DeepSeek-R1 [13], a reasoning model exhibiting performance comparable to OpenAI’s O1 model, available for open research [9]. DeepSeek-R1 is trained through large-scale reinforcement learning (RL) without supervised fine-tuning, demonstrating powerful reasoning behaviors [13].
@@
We present five examples to assess the model performance. The first example in this study features a conjugate gradient solver that incorporates matrix and vector operations. The second example is a parallel 1D stencil-based heat equation solver [7], chosen for its relevance in testing parallel computational workflows and its widespread use in scientific computing. The third example is parallel matrix multiplication, a fundamental operation in scientific computing and machine learning, often accelerated using HPC systems to enhance computational speed. The fourth example is parallel double-precision general matrix multiply (DGEMM), a widely used operation in numerical linear algebra that is essential for many scientific and engineering applications, testing the system’s computational efficiency for large-scale matrix operations. The final example is the STREAM triad operation, a key benchmark for testing memory bandwidth and data movement between CPU and memory, commonly used to evaluate the memory subsystem’s performance in HPC systems. We summarize the main contributions of this paper as follows:
@@ 2 Related Work
LLMs have been widely used across various domains, including natural language processing (NLP) and visualization. However, their application in analyzing and optimizing HPC tasks remains challenging. In Table 1, we present the related works that use LLMs for HPC tasks. LLM4HPC [4] represents the first effort to adapt LLMs specifically for the HPC domain. The LLM4HPC framework is specifically designed for HPC, showing success in code similarity analysis, parallelism detection, and OpenMP question-answering tasks. LLM4VV [20] is a fine-tuned model that leverages the capabilities of GPT-4 and CodeLlama (based on Llama-2) to successfully generate OpenACC directives. HPC-GPT introduced by Ding et al. [8], was built on LLaMa-13B. HPC-GPT has been successfully applied to managing AI models and datasets, as well as detecting data races. Kadosh et al. [15] introduced the domain-specific Tokompiler LLM, which outperforms a GPT-3-based model in code completion and semantics for Fortran, C, and C++ code. Nichols et al. [21] present HPC-Coder. Their work demonstrated varying success in code completion, including handling OpenMP pragmas and MPI calls. Lei et al. [16] introduced a dataset designed for fine-tuning models on OpenMP Fortran and C++ code translation. Their fine-tuned model yielded more accurate results compared to GPT-4. Chen et al. [3] provide an insightful overview of the challenges and opportunities at the intersection of LLMs and HPC, extending beyond code generation. Notably, the works by Godoy et al. [10] and Valero-Lara et al. [25], which evaluate HPC kernel code generation and results for LLaMa-2, are among the first to apply LLM-based code generation to the domain of HPC software development. Additionally, Godoy et al. [11] apply LLM capabilities of GPT-3 targeting HPC kernels for code generation, and auto-parallelization of serial code in C++, Fortran, Python and Julia. Yin et al. [26] proposed chatHPC, a chatbot for HPC question answering and script generation.
@@ 3 Methodology
Example 3: In this example, we evaluate DeepSeek’s ability to generate parallel code for matrix multiplication, a fundamental operation in numerical simulations and machine learning. Matrix multiplication is a critical task in scientific computing, particularly in HPC, where large matrices are multiplied to generate results in parallel. The task involves writing efficient parallel code to split matrix multiplication across multiple threads or processors, optimizing for memory usage and computational speed. The generated code is supposed to compute the following:
@@
Example 4: The fourth example is more general and focuses on DGEMM, a well known operation in scientific computing and HPC. The objective here is to analyze the ability of DeepSeek to generate parallel code for performing DGEMM on large-scale matrices, with a particular focus on optimizing for performance using parallel computing techniques. The generated code is supposed to compute the following:
@@ 4.5 STREAM
Figure 1: Two-dimensional classification using the estimated schedule effort of the COCOMO model (easy vs difficult) and the quality of the code using compilation, execution, and correctness (poor vs good).
@@ 5 Performance evaluation
For the performance evaluation, the following versions were used: g++/gfortran 13/14, Julia 1.11.3 and Python 3.11. Recall that we just fixed compilation and runtime errors in the code, but did not improve the parallel implementations. Figure 2(a) shows the scaling for the generated codes on AMD EPYC 7763 (x86) from one thread to 64 cores. We used $10,000,000$ nodes and changed the length of the domain $L$ to $100,000$. The C++ code and Python code scaled with the number of cores. The Fortran and Julia code showed some speedup from a single core to five cores. After that, the code did not benefit from additional cores. Figure 2(b) shows the scaling from one core to 64 cores on Intel Xeon Platinum 8358 (x86). The matrix size was $000\times 10,$ 000 × 10 , 000 with $100,000,000$ elements. Here, only the Julia, code scaled. Python, C++  and Fortran showed strange behavior. Figure 2(c) shows the DGEMM benchmark for matrices with 512, 1024, and 2048 rows and columns and a block width of 64 on Arm A64FX. These numbers were generated by the LLM. The Fortran code was very slow and the GFLOP/s were in the range of 0.02. The Python code reported around the same values for all sizes. The C++ and Julia code had increasing numbers. However, the Julia codes reported GFLOP/s for a BLAS implementation and these values were hundreds times bigger. Thus, all of the codes had a very poor performance. Figure 2(d) shows the performance for stream triad for $10^{4}$, $10^{5}$, and $10^{6}$ elements on ARM Grace Grace using 72 cores. Here, all code increased performance with the array size. For the Fortran code, we used opencoarrays with openmpi 4.1 as the coarray implementation.
@@
Figure 2: Performance measurements: (2(a)) parallel heat equation solver on x86, (2(b)) parallel matrix multiplication on RISC-V, (2(c)) DGEMM on Arm A64FX, and (2(d)) stream triad on Arm Grace Hopper.
@@ 6 Discussion and Conclusion
One important observation is that LLMs struggle to write scalable codes. In first two examples, both the C++ and Python codes showed consistent scaling with the increase in the number of cores, while the Fortran and Julia implementations showed limited scalability. This is consistent with prior research on HPC applications, which highlights the challenges of parallelism in certain programming languages. In example three, both C++ and Fortran scaled, but the Julia code could not be executed on RISC-V CPUs due to Julia not supporting this device. The Stream Triad clearly improved in performance with increasing array sizes for the all generated codes. Using COCOMO analysis, the C++  code required more effort in the conjugate gradient and matrix multiplication examples. While the effort was more balanced between all the languages for the heat equation solver and STREAM Triad tasks.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have rapidly advanced, potentially approaching Artificial General Intelligence (AGI). Trained on vast textual data, LLMs have shown remarkable capabilities in natural language processing and visualization tasks [1, 15, 25]. This trend has accelerated with the release of the DeepSeek models by the Chinese company DeepSeek [7]. Their DeepSeek-V3 [20, 17] was trained in two months at a cost of \$5.6 million, significantly cheaper than similar models. In contrast, the development cost of ChatGPT-4 ranged between \$41 million and \$78 million [3]. On January 20th 2025, they launched DeepSeek-R1 [16], a reasoning model exhibiting performance comparable to OpenAI’s O1 model, available for open research [12]. DeepSeek-R1 is trained through large-scale reinforcement learning (RL) without supervised fine-tuning, demonstrating powerful reasoning behaviors [16].
@@
We present five examples to assess the performance of the model. The first example in this study features a conjugate gradient solver that incorporates matrix and vector operations. The second example is a parallel 1D stencil-based heat equation solver [10], chosen for its relevance in testing parallel computational workflows and its widespread use in scientific computing. The third example is parallel matrix multiplication, a fundamental operation in scientific computing and machine learning, often accelerated by using HPC systems to enhance computational speed. The fourth example is parallel double-precision general matrix multiply (DGEMM), a widely used operation in numerical linear algebra that is essential for many scientific and engineering applications, testing the system’s computational efficiency for large-scale matrix operations. The final example is the STREAM triad operation, a key benchmark for testing memory bandwidth and data movement between CPU and memory, commonly used to evaluate the memory subsystem’s performance in HPC systems. We summarize the main contributions of this paper as follows:
@@ 2 Related Work
LLMs have been widely used across various domains, including natural language processing (NLP) and visualization. Multiple efforts have also been made to benchmark LLMs for tasks such as code generation [9, 8, 6]. However, their application in analyzing and optimizing HPC tasks remains challenging. In Table 1, we present the related works that use LLMs for HPC tasks. LLM4HPC [5] represents the first effort to adapt an LLM specifically to the HPC domain. The LLM4HPC framework is specifically designed for HPC, showing success in code similarity analysis, parallelism detection, and OpenMP question-answering tasks. LLM4VV [23] is a fine-tuned model that uses the capabilities of GPT-4 and CodeLlama (based on Llama-2) to successfully generate OpenACC directives. HPC-GPT introduced by Ding et al. [11], was built on LLaMa-13B. HPC-GPT has been successfully applied to managing AI models and datasets, as well as detecting data races. Kadosh et al. [18] introduced the domain-specific Tokompiler LLM, which outperforms a GPT-3-based model in code completion and semantics for Fortran, C, and C++ code. Nichols et al. [24] present HPC-Coder. Their work demonstrated varying success in code completion, including handling OpenMP pragmas and MPI calls. Lei et al. [19] introduced a dataset designed for fine-tuning models on OpenMP Fortran and C++ code translation. Their fine-tuned model yielded more accurate results compared to GPT-4. Chen et al. [4] provide an insightful overview of the challenges and opportunities at the intersection of LLMs and HPC, extending beyond code generation. Notably, the works by Godoy et al. [13] and Valero-Lara et al. [28], which evaluate HPC kernel code generation and results for LLaMa-2, are among the first to apply LLM-based code generation to the domain of HPC software development. Additionally, Godoy et al. [14] apply LLM capabilities of GPT-3 targeting HPC kernels for code generation, and auto-parallelization of serial code in C++, Fortran, Python and Julia. Yin et al. [29] proposed chatHPC, a chatbot for HPC question answering and script generation.
@@ 3 Methodology
Example 3: In this example, we evaluate DeepSeek’s ability to generate parallel code for matrix multiplication, a general benchmarking task used to evaluate the basic performance and scalability of parallel code. It is also a fundamental operation in numerical simulations and machine learning. The task involves writing efficient parallel code that runs across multiple threads or processors, optimizing for memory usage and computational speed. The generated code is supposed to compute the following:
@@
Example 4: The fourth example is more general and focuses on DGEMM, which targets a more specialized application within the HPC domain. The objective here is to analyze the ability of DeepSeek to generate parallel code for performing DGEMM on large-scale matrices, with a particular focus on optimizing for performance using parallel computing techniques. The generated code is supposed to compute the following:
@@
While Example 3 is useful for assessing general parallel performance, DGEMM provides a deeper evaluation of optimized performance for large-scale, high-precision operations.
@@ 4.5 STREAM
Figure 1: Two-dimensional classification using the estimated schedule effort of the COCOMO model (easy vs difficult) and the quality of the code using compilation, execution, and correctness (poor vs good). The blue values show the results from ChatGPT [9] and the black values the results for Deep Seek. The Python and Julia data points are tagged with the common file endings py and jl, respectively.
@@ 5 Performance evaluation
For the performance evaluation, the following versions were used: g++/gfortran 13/14, Julia 1.11.3 and Python 3.11. Recall that we just fixed compilation and runtime errors in the code, but did not improve the parallel implementations. Figure 2(a) shows the scaling for the generated codes on AMD EPYC 7763 (x86) from one thread to 64 cores. We used $10,000,000$ nodes and changed the length of the domain $L$ to $100,000$. The C++ code and Python code scaled with the number of cores. The Fortran and Julia code showed some speedup from a single core to five cores. After that, the code did not benefit from additional cores. Figure 2(b) shows the scaling from one core to 64 cores on Intel Xeon Platinum 8358 (x86). The matrix size was $000\times 10,$ 000 × 10 , 000 with $100,000,000$ elements. Here, only the Julia, code scaled. Python, C++  and Fortran showed strange behavior. Figure 2(c) shows the DGEMM benchmark for matrices with 512, 1024, and 2048 rows and columns and a block width of 64 on Arm A64FX. These numbers were generated by the LLM. The Fortran code was very slow and the GFLOP/s were in the range of 0.02. The Python code reported around the same values for all sizes. The C++ and Julia code had increasing numbers. However, the Julia codes reported GFLOP/s for a BLAS implementation and these values were hundreds times bigger. Thus, all of the codes had a very poor performance. Figure 2(d) shows the performance for stream triad for $10^{4}$, $10^{5}$, and $10^{6}$ elements on ARM Grace Grace using 72 cores. Here, all code increased performance with the array size. For the Fortran code, we used opencoarrays with openmpi 4.1 as the coarray implementation. Table 5 shows the peak performance for the parallel heat equation and matrix multiplication.
@@
Figure 2: Performance measurements: (2(a)) parallel heat equation solver on x86-AMD, (2(b)) parallel matrix multiplication on x86-Intel, (2(c)) DGEMM on Arm A64FX, and (2(d)) stream triad on Arm Grace Hopper.
Table 5: Peak performance for the parallel heat equation and parallel matrix multiplication as points processed per second and matrix entries processed per second, respectively. The number in parentheses are the number of cores at the peak performance.
Example C++ Fortran Python Julia
Heat equation 64 790 3386479033864\,790\,33864 790 338 (64) 604 750604750604\,750604 750 (15) 8 127 13781271378\,127\,1378 127 137 (64) 166 899166899166\,899166 899 (15)
Matrix multiplication 192 661192661192\,661192 661 (60) 7 502 60775026077\,502\,6077 502 607 (25) 169 678169678169\,678169 678 (60) 1 225 91912259191\,225\,9191 225 919 (55)
@@ 6 Discussion and Conclusion
One important observation is that LLMs struggle to write scalable codes. In first two examples, both the C++ and Python codes showed consistent scaling with the increase in the number of cores, while the Fortran and Julia implementations showed limited scalability. This is consistent with prior research on HPC applications, which highlights the challenges of parallelism in certain programming languages. In example three, both Python and Fortran did not scale. The Stream Triad clearly improved in performance with increasing array sizes for the all generated codes. Using COCOMO analysis, the C++  code required more effort in the conjugate gradient and matrix multiplication examples. While the effort was more balanced between all the languages for the heat equation solver and STREAM Triad tasks.